# Grid-based Multi-Robot Coverage Path Planning

### v1

@@ 20                   $O_{d}\leftarrow O_{d}\cup\{o_{d}^{\text{(v)}}=(i,\{v\})\}$
Properties: After any valid operator is executed, (1) every $G_{i}\in\mathcal{G}$ remains connected since only deduplicate operators can remove edges but a valid deduplicate guarantees that $G_{i}$ remains connected, and (2) the property of the union of the vertices of all subgraphs being the vertex set $V$ is not affected since no vertex coverage is lost. Therefore, with Theorem 3, we know that the coverage paths in an LS-MCPP solution $\Pi$ jointly cover all vertices of $G$ if $\bigcup_{i\in I}V_{i}=V$.
V-C Heuristics for Operator Sampling
To better explore the solution neighborhood, LS-MCPP should carefully determine which operator to sample from the selected pool (line 2 of Alg. 2). To this end, we propose three heuristic functions tailored to the three types of operators to evaluate their potential in guiding the neighborhood search and improving the solution. An operator with a larger heuristic value results in a higher probability of being sampled. For the grow operator, its heuristic function is determined by two considerations: (1) prioritizing growing light subgraphs with smaller path costs using ESTC and (2) prioritizing the edges whose two endpoints have less duplication. Formally, the heuristic value $h(o_{g})$ for a grow operator $o_{g}(i,V_{o_{g}})$ is defined as $h(o_{g})=-k\cdot c(\pi_{i})-\frac{1}{|V_{o_{g}}|}\cdot\sum_{v\in V_{o_{g}}}n_{v}$. Here, $k\geq n_{z}$ holds for any $z\in V$, prioritizing consideration (1) over consideration (2). For deduplicate operator $o_{d}(i,V_{o_{d}})$, the heuristic function is the opposite of the one for grow operators: $h(o_{d})=k\cdot c(\pi_{i})+\frac{1}{|V_{o_{d}}|}\cdot\sum_{v\in V_{o_{d}}}n_{v}$. For exchange operator $o_{e}(i,j,V_{o_{e}})$, its heuristic function is defined as $h(o_{e})=c(\pi_{j})-c(\pi_{i})$ to prioritize two subgraphs with a larger difference between their path costs.
@@ VI Deconflicting MCPP Solutions
In Sec. V, the MCPP solution produced by LS-MCPP does not inherently guarantee that the robots will be conflict-free during their operations in practice. We observe that the robots would be conflict-free in most regions of the graph, as they tend to naturally disperse to cover different regions in an optimized LS-MCPP solution to minimize the overall makespan. Despite this, inter-robot collisions remain a significant challenge in certain congested areas, such as the region of root vertices of the robots when they are clustered or the narrow corridors that only allow one robot to pass through at a time. To tackle this challenge, we leverage techniques from MAPF research, which focuses on finding a set of collision-free paths for a group of robots to navigate across a grid graph. Following the common practice for solving MAPF problems, we develop a two-level hierarchical planner to deconflict an MCPP solution, where the high-level planner determines how the conflicts between the robots are resolved, and the low-level planner focuses on single-robot planning by considering other robots as spatial-temporal obstacles.
VI-A Preliminary
We aim to construct a set $\mathcal{P}=\{P_{i}\}_{i=1}^{k}$ of $k$ conflict-free plans from an input solution $\Pi=\{\pi_{i}\}_{i=1}^{k}$ for the MCPP instance $(G,I,R)$. Each plan $P_{i}=(\mathbf{x}_{1},\mathbf{x}_{2},...,\mathbf{x}_{|P_{i}|})$ is an ordered sequence of states subject to the order of the vertices in $\pi_{i}=(v_{1},v_{2},...,v_{|\pi_{i}|})$. The time-embedded state $\mathbf{x}=(v,t)$ represents the state of a robot arriving vertex $v\in V$ at time $t\geq 0$, and the initial state of each robot $i$ is set as $\mathbf{x}_{1}=(r_{i},0)\in P_{i}$ where $r_{i}\in R$. The time cost $c(P_{i})$ of robot $i$ is the associated time of the last state $\mathbf{x}_{|P_{i}|}$. For convenience, we use $P[j]$, $V(P)\subseteq V$, and $E(P)$ to denote the $j^{\text{th}}$ state $\mathbf{x}_{i}$ of any plan $P$, the set of unique vertices visited by $P$, and the list of (potentially duplicated) edges used by $P$, respectively. For any adjacent vertices $v_{j},v_{j+1}\in\pi_{i}$, a subsequence of plan $P_{i}$ is constructed to transit from state $(v_{j},\cdot)$ to $(v_{j+1},\cdot)$ without any conflict (see Definition 5). To transit from a state $\mathbf{x}=(v,t)$ to another state $\mathbf{x}^{\prime}=(v^{\prime},t^{\prime})$ in a plan, the robot can either wait at vertex $v$ for a time period $t^{\prime}-t>0$ when $v=v^{\prime}$, or move from vertex $v$ to vertex $v^{\prime}\neq v$ via edge $e=(v,v^{\prime})\in E$ at the time cost of $t^{\prime}-t=w_{e}>0$.
@@ VI-A Preliminary
Consider any two plans $P_{a}$ and $P_{b}$ where each state of $P_{a}$ is $\mathbf{x}^{a}_{i}=(v_{i}^{a},t_{i}^{a})$ for $i=1,2,...,|P_{a}|$ and each state of $P_{b}$ is $\mathbf{x}^{b}_{j}=(v_{j}^{b},t_{j}^{b})$ for $j=1,2,...,|P_{b}|$. A conflict on vertex $v\in V$ occurs if there exist $\mathbf{x}^{a}_{i}\in P_{a}$ and $\mathbf{x}^{b}_{j}\in P_{b}$ such that $v^{a}_{i}=v^{b}_{j}=v$ and the two time intervals $[t_{i-1}^{a},t_{i+1}^{a})$ and $[t_{j-1}^{b},t_{j+1}^{b})$ intersects with each other.
Note that two dummy states $\mathbf{x}_{0}=(v_{1},0)$ and $\mathbf{x}_{|P|+1}=(v_{|P|},+\infty)$ are padded before and after each plan $P$ to check if there is a conflict at $\mathbf{x}_{1}$ or $\mathbf{x}_{|P|}$, since the robot associated with $P$ will stay at $v_{1}$ before the task and at $v_{|P|}$ after the task. Our definition of conflict is more conservative than that in most MAPF algorithms. Specifically, we assume that a vertex is blocked not only when it is occupied by a robot, as is typical in the MAPF literature [sharon2015conflict], but also when a robot is moving out of it. By checking for time intervals during which other robots occupy a vertex, we can implicitly identify all conflict types outlined in the literature [stern2019multi], including vertex conflicts, edge conflicts, following conflicts, and swapping conflicts. These conflicts are typically defined for unweighted graphs, where time is discretized into uniform intervals for the robots, allowing their actions to be aligned step by step.
Theorem 7.
It is NP-hard to find an optimal deconflicted MCPP solution with a minimized makespan.
{proof}
The OMG-MAPF problem (see Sec. II-B) was proved to be NP-hard with a makespan minimization objective [mouratidis2024fools], which can be viewed as a special case of our problem that operates solely on unweighted graphs (i.e., $w_{e}=1$ for each $e\in E$). Therefore, it is at least NP-hard to find an optimal deconflicted MCPP solution with a minimized makespan.
@@
Instead of first solving MCPP and then deconflicting the resulting MCPP solution as proposed, one could also deconflict the MCPP solution at the end of each iteration of LS-MCPP (Alg. 2), which holds the potential to outperform the proposed sequential approach. However, it incurs significantly higher time complexity due to the computational demands of the deconflicting procedure (see Theorem 7 and Tab. II).
@@ VI-B The High-Level Planner
The high-level planner determines how the conflicts defined in Definition 5 are resolved, which builds a search tree by growing child nodes whenever a conflict is detected. At each node of the search tree, a low-level planner is used to compute a plan for each robot with the help of a reservation table to avoid the conflicts that are reserved in it. In short, the reservation table is a dictionary that stores the items with the vertex as key and a set of disjoint time intervals as value.
Figure 11: The high-level Prioritized-Based Search (PBS) planner.
In consideration of time complexity as demonstrated in Theorem 7, we use Priority-based Search (PBS) [ma2019searching] as the high-level planner. PBS maintains a search tree of nodes, where each node $n$ is associated with a set $\mathbf{o}_{n}$ of pairwise partial orders and a set $\mathcal{P}_{n}$ of plans. As illustrated in Fig. 11, PBS resolves the pairwise conflicts between robot $i$ and $j$ at each node $n$ by splitting it into two child nodes, $a$ and $b$, one with the order $i\prec j$ to $\mathbf{o}_{a}$ added and the other with order $j\prec i$ added to $\mathbf{o}_{b}$. For each child node (e.g., node $a$), the corresponding $\mathcal{P}_{a}$ is updated by invoking a low-level planner to update the plan of each robot affected by the addition of the order $j\prec i$ into $\mathbf{o}_{a}$. Each plan is computed with all the plan states from all robots with higher priority reserved in the reservation table. The search tree grows in a depth-first search (DFS) fashion until it reaches a leaf node $n$ where all conflicts in set $\mathcal{P}_{n}$ are resolved using $\mathbf{o}_{n}$. The following remark explains that compared with CBS, PBS is more time efficient due to a much smaller DFS tree, since each search node resolves all the conflicts between the two robots.
@@
Alternatively, one can employ Conflict-based Search (CBS) [sharon2015conflict] as the high-level planner. CBS addresses pairwise conflicts between robots $i$ and $j$ in each search node by splitting it into two child nodes. In one child node, the plan for robot $i$ is updated by invoking the low-level planner with the conflicted state added to its reservation table; in the other child node, robot $j$’s plan is updated with the conflicted state recorded in its reservation table. As the search tree expands, CBS systematically resolves all possible conflicts. CBS-like algorithms are proven to be optimal for various MAPF variants when the corresponding low-level single-robot planner is optimal. However, this approach often requires longer runtimes to explore a larger number of nodes and construct the search tree necessary to find the optimal solution. This is particularly true in our multi-goal setting of MAPF, where the optimal low-level planner has significantly higher time complexity. According to [mouratidis2024fools], OMG-MAPF, a special case of our deconflicting problem in Sec. VI-A where the edges are unweighted was solved with CBS as the high-level planner, but the method could only scale to sequences of up to 5 goals on relatively small graphs.
@@ VI-C The Low-level Planner
In the lower-level planner, our goal is to compute a partially conflict-free plan $P_{i}$ for robot $i$ to sequentially visit the ordered set of vertices along its path $\pi_{i}\in\Pi$. The plan only avoids conflicts with higher-priority plans by reserving their states in the reservation table. To accommodate planning in graph $G$ with nonuniform edge weights, we adopt the idea from Safe Interval Path Planning (SIPP) [phillips2011sipp] for single-goal shortest path planning. In essence, SIPP expands upon the heuristic search concept of space-time A [hart1968formal] with the only difference on how the states are stored in the reservation table. Given a state $\mathbf{x}=(v,t)$ to be recorded in the reservation table, space-time A records the discrete time $t$ for vertex $v$, whereas SIPP records the time interval $[t^{\prime},t)$ for vertex $v$ where $t^{\prime}$ is the arrival time of the preceding state $\mathbf{x}^{\prime}=(v^{\prime},t^{\prime})$ of $\mathbf{x}$. Inspired by [mouratidis2024fools], we develop three approaches for the low-level planner upon SIPP, namely chaining approach, holistic approach, and a novel adaptive approach. Before the low-level planning, we do a preprocessing step on each $\pi_{i}\in\Pi$ to remove any $v\in R\setminus\{r_{i}\}$ from $\pi_{i}$. This is crucial to make the low-level planner more robust and avoid a doomed failure when a root vertex of another robot is an intermediate goal in the goal sequence.
Figure 12: A PBS node with partial order $2\prec 1$. Left: first conflict found at vertex $z$ between state $(z,3)$ of robot $1$ and state $(z,2)$ of robot $2$, where the edge weights are annotated along the edge, and the two paths $\pi_{1}$ and $\pi_{2}$ are colored in red and blue, respectively. Right: plans $P_{1}$ and $P_{2}$ of the two robots and the reservation table $\text{rt}_{1}$ of robot $1$. The chaining approach cannot resolve the conflict using SIPP, as both $u$ and $z$ are unreachable once robot $1$ reaches state $\mathbf{x}=(v,2)$. The holistic and adaptive approaches can resolve the conflict using ML-SIPP when reaching state $\mathbf{x}$ by backtracking to states expanded from $(u,0)$.
Chaining Approach: Given a path $\pi=(v_{1},v_{2},...,v_{|\pi|})$ for any robot, the chaining approach considers each pair of consecutive vertices of $\pi$ as a single-robot path planning problem from a goal state $\mathbf{x}_{i}^{(\pi)}=(v_{i},\cdot)$ to the next goal state $\mathbf{x}_{i+1}^{(\pi)}=(v_{i+1},\cdot)$. We then use SIPP with the admissible true distance heuristic [silver2005cooperative], denoted as $h^{T}(\mathbf{x},\mathbf{x}^{(\pi)})$. It requires an auxiliary backward search from the goal state $\mathbf{x}^{(\pi)}$ to compute the minimum cost to any state $\mathbf{x}$ ignoring any dynamic obstacles in the time dimension. Note that reaching a goal state $\mathbf{x}^{(\pi)}$ in SIPP is determined by the vertex regardless of its arrival time. By concatenating the resulting consecutive plans one after the other, we obtain the final plan $P_{i}$ for the robot $i$. According to [mouratidis2024fools], the chaining approach can be shown to be suboptimal and also not complete. Such property persists when we directly replace the single-robot planner of space-time A with SIPP, as demonstrated in the example of Fig. 12. Intuitively speaking, its suboptimality and incompleteness result from its restricted search space.
Holistic Approach: The holistic approach borrows the similar idea from Multi-Label A [grenouilleau2019multi, zhong2022optimal], originally for multi-goal MAPF variants in unweighted graphs. By incorporating a label dimension into the robot state, we adopt Multi-Label A for our weighted graph settings with SIPP, namely Multi-Label Safe-Interval Path Planning (ML-SIPP). Specifically speaking, for each vertex $v_{i}$ in the path $\pi=(v_{1},v_{2},...,v_{|\pi|})$, we define its corresponding goal state as $\mathbf{x}_{i}^{(\pi)}=(v_{i},\cdot,l_{i})$ with label $l_{i}=i$ indicating $v_{i}$ as its current goal vertex to reach. Here, the arrival time can be arbitrary, consistent with the goal-reaching condition on only the vertex dimension as in the chaining approach. The transition between any two consecutive states $\mathbf{x}_{j}$ and $\mathbf{x}_{j+1}$ in a plan is thereby updated for the label dimension, regardless of the inducing actions: if $\mathbf{x}_{j}$ is a goal state with label $l$, the label of $\mathbf{x}_{j+1}$ is set as $l+1$; otherwise, the label inherits the label $l$ of $\mathbf{x}_{j}$. To further tighten the true distance heuristics $h^{T}$ in SIPP, the same concatenated heuristics from Multi-Label A for ML-SIPP:
$h(\mathbf{x},\mathbf{x}_{i})=h^{T}(\mathbf{x},\mathbf{x}_{i}^{(\pi)})+\sum_{j=i}^{|\pi|-1}h^{T}(\mathbf{x}_{j}^{(\pi)},\mathbf{x}_{j+1}^{(\pi)}),$ (5)
where $\mathbf{x}_{i}^{(\pi)}$ is the current goal state to reach updated throughout the search procedure. One can easily verify that $h$ is admissible as $h^{T}$ is admissible, with which we can prove the holistic approach is an optimal and complete low-level planner.
Adaptive Approach: So far we have introduced two approaches of the low-level planner: the chaining approach is fast yet incomplete, whereas the holistic approach is complete and optimal yet has large time and space complexity with the additional label dimension. To improve the practicability of both approaches, we propose a novel adaptive low-level planner in Alg. 5 combining their benefits. Instead of running ML-SIPP on the entire sequence of goals as in the holistic approach, our adaptive approach calls ML-SIPP on a limited window of the goal sequences that traces back to previous goals whenever a sequence containing only the current goal is unsolvable by ML-SIPP [Lines 5-5]. The adaptive windows size is restricted by the input max backtracking size parameter $b_{\text{max}}\in\mathbb{Z}^{+}$. When ML-SIPP finds a valid plan for the sequence containing only the current goal, it proceeds to the next goal and resets the current backtracking window $b$ [Lines 5-5]. Note that ML-SIPP falls back to SIPP when the sequence $S_{\text{goal}}$ of goals states contains only one state. Our adaptive approach is complete since it returns the ML-SIPP solution to sequentially visit the goals in $\pi$ in the worst-case scenario [Lines 5-5]. In practice, we find that setting $b_{\text{max}}$ to a small integer is sufficient enough to resolve all conflicts even with a limit on the number of expanded nodes, as we will see in Sec. VII.
Input: path $\pi=(v_{1},v_{2},...,v_{|\pi|})$ for robot $i$

### v2

@@ 20                   $O_{d}\leftarrow O_{d}\cup\{o_{d}^{\text{(v)}}=(i,\{v\})\}$
Properties: After any valid operator is executed, (1) every $G_{i}\in\mathcal{G}$ remains connected since only deduplicate operators can remove edges but a valid deduplicate operator guarantees that $G_{i}$ remains connected, and (2) the property of the union of the vertices of all subgraphs being the vertex set $V$ is not affected since no vertex coverage is lost. Therefore, an LS-MCPP solution ensures complete coverage due to Theorem 3.
V-D Heuristics for Operator Sampling
To better explore the solution neighborhood, LS-MCPP should carefully determine which operator to sample from the selected pool (line 2 of Alg. 2). We propose three heuristic functions tailored to the three types of operators to evaluate their potential in guiding the neighborhood search and improving the solution. An operator with a larger heuristic value results in a higher probability of being sampled. The heuristic function for grow operators has two considerations: (1) prioritizing growing light subgraphs with smaller path costs and (2) prioritizing covering vertices with less duplication. Formally, the heuristic value $h(o_{g})$ for a grow operator $o_{g}(i,V_{o_{g}})$ is defined as $h(o_{g})=-k\cdot c(\pi_{i})-\frac{1}{|V_{o_{g}}|}\cdot\sum_{v\in V_{o_{g}}}n_{v}$. Note that $k\geq n_{v}$ holds for any $v\in V$, thereby prioritizing consideration (1) over consideration (2). The heuristic function for deduplicate operators $o_{d}(i,V_{o_{d}})$ is defined as the exact opposite of the one for grow operators: $h(o_{d})=k\cdot c(\pi_{i})+\frac{1}{|V_{o_{d}}|}\cdot\sum_{v\in V_{o_{d}}}n_{v}$. The heuristic function for exchange operators $o_{e}(i,j,V_{o_{e}})$ is defined as $h(o_{e})=c(\pi_{j})-c(\pi_{i})$ to prioritize pairs of subgraphs with larger differences in their path costs.
@@ VI Deconflicting MCPP Solutions
In this section, we address inter-robot conflicts in coverage paths produced by LS-MCPP or other grid-based MCPP methods. Our MCPP formulation (Problem 1) does not inherently prevent collisions among robots. In an MCPP solution with a small makespan, robots typically cover different regions and are distributed across the grid, which lowers the likelihood of conflicts. However, certain regions remain prone to congestion and conflicts, particularly around clustered root vertices or narrow corridors that only allow one robot to pass at a time.
To resolve these potential conflicts, we formulate the deconflicting task as a variant of MAPF, leveraging MAPF techniques to ensure each robot follows its designated coverage path while avoiding collisions with others. Specifically, we develop a two-level hierarchical planner, following standard MAPF practices, to effectively deconflict MCPP solutions: The high-level planner employs Priority-Based Search (PBS) [60] to assign priorities among pairs of conflicting robots; The low-level planner is a novel algorithm that plans a continuous-time path for each robot that avoids conflicting with any higher-priority robots. Our key extensions from standard MAPF lie in 1) our treatment of non-uniform edge weights, which translate to variable time costs for robot actions and 2) an adaptive approach of the low-level planner for time-efficient single-robot path planning with multiple goals.
VI-A Formulation of Deconflicted MCPP
Given an input solution $\Pi=\{\pi_{i}\}_{i=1}^{k}$ for the MCPP instance $(G,I,R)$, we define the Deconflicted MCPP problem, which aims to construct a set $\mathcal{T}=\{\tau_{i}\}_{i=1}^{k}$ of $k$ conflict-free trajectories. A trajectory $\tau_{i}$ for robot $i$ is an ordered sequence $(\mathbf{x}_{1},\ldots,\mathbf{x}_{|\tau_{i}|})$ of time-embedded states $\mathbf{x}=(v,t)$ indicating that the robot stays at vertex $v\in V$ at time $t\geq 0$. Given path $\pi_{i}=(v_{1},\ldots,v_{|\pi_{i}|})$ where $v_{1}=v_{|\pi_{i}|}=r_{i}\in R$, trajectory $\tau_{i}$ must start from state $\mathbf{x}_{1}=(v_{1},0)$, visit all vertices in the specified order of $\pi_{i}$, and end with state $\mathbf{x}_{|\tau_{i}|}=(v_{|\pi_{i}|},\cdot)$. Formally, there exist indices $a_{1}=1<a_{2}<\ldots<a_{|\pi_{i}|}=|\tau_{i}|$ such that $\mathbf{x}_{a_{j}}=(v_{j},\cdot)$ for all $j=1,\ldots,|\pi_{i}|$. To transit from state $(u,t)$ to the next state $(v,t^{\prime})$ in a trajectory, the robot performs a wait-and-move action, where it waits at $v$ for $t^{\prime}-w_{e}-t\geq 0$ and then moves from $u$ to $v$ along edge $e=(u,v)\in E$ with a time cost of $w_{e}>0$. Let $\tau[j]$, $\mathbf{x}.v$ and $\mathbf{x}.t$ denote the $j$-th state in $\tau$, the vertex and time in each state $\mathbf{x}$, respectively. The quality of a Deconflicted MCPP solution $\mathcal{T}$ is evaluated using the makespan metric, denoted by $\theta(\mathcal{T})=\max\{\tau_{1}[|\tau_{1}|].t,\ldots,\tau_{k}[|\tau_{k}|].t\}$, which is consistent with that used for MCPP.
@@ VI-A Formulation of Deconflicted MCPP
A conflict between trajectory $\tau$ with states $\mathbf{x}_{j}$ for $j=1,\ldots,|\tau|$ and another trajectory $\tau^{\prime}$ with states $\mathbf{x}^{\prime}_{j^{\prime}}$ for $j^{\prime}=1,\ldots,|\tau^{\prime}|$ occurs if there exist states $\mathbf{x}_{j}$ and $\mathbf{x}^{\prime}_{j^{\prime}}$ such that their vertices $\mathbf{x}_{j}.v=\mathbf{x}^{\prime}_{j^{\prime}}.v$ and the two time intervals $(\mathbf{x}_{j-1}.t,\mathbf{x}_{j+1}.t)$ and $(\mathbf{x}^{\prime}_{j^{\prime}-1}.t,\mathbf{x}^{\prime}_{j^{\prime}+1}.t)$ intersect. To ensure conflict checks are valid for all $j$, we assume two dummy bounds $\mathbf{x}_{0}.t=0$ and $\mathbf{x}_{|\tau|+1}.t=+\infty$.
In our definition of conflicts, a vertex is considered occupied by a robot not only when the robot stays at it (as in typical MAPF frameworks [61]) but also when a robot is moving into or out of it. This conservative definition subsumes all types of conflicts outlined in MAPF literature [17], including vertex conflicts, edge conflicts, following conflicts, and swapping conflicts. This generalized definition accommodates continuous time and non-uniform edge weights, unlike traditional MAPF definitions that assume discretized, uniform time steps.
@@
Instead of first solving MCPP and then deconflicting the resulting paths as a post-processing procedure, one could integrate the deconfliction at the end of each LS-MCPP iteration (Alg. 2), potentially improving performance beyond our proposed sequential approach. However, such an approach incurs significantly higher runtime due to the computational demands of repeated runs of the deconfliction (see Tab. II).
@@ VI-B The High-Level Planner
The high-level planner employs Priority-Based Search (PBS) [60] to resolve conflicts defined in Definition 1. PBS searches a priority tree where each node $N$ stores a unique priority ordering given by a set $\bm{\boldsymbol{\prec}}_{N}$ of ordered pairs of robots and a set $N.\mathcal{T}$ of $k$ trajectories that respect the prioritized planning scheme specified by $\bm{\boldsymbol{\prec}}_{N}$. The root node contains an empty priority set and a set of potentially conflicting trajectories. When PBS expands a node $N$, it checks $N.\mathcal{T}$ for conflicts. If no conflicts are found, then $N$ is a goal node, and $N.\mathcal{P}$ is the solution. If conflicts exist between the trajectories of robots $i$ and $j$, PBS generates two child nodes, $N_{1}$ and $N_{2}$, and adds the pair $i\prec j$ (robot $i$ has a higher priority than robot $j$) to $N_{1}$ and $j\prec i$ to $N_{2}$. For child node $N_{1}$, PBS invokes a low-level planner to update the trajectories of robot $j$ and other robots $j^{\prime}$ with $j\prec_{N_{1}}j^{\prime}$, ensuring $N_{1}.\mathcal{T}$ respects $\bm{\boldsymbol{\prec}}_{N_{1}}$. PBS processes $N_{2}$ symmetrically. The low-level planner computes a trajectory for each robot that avoids conflicts with higher-priority trajectories. PBS performs Depth-First Search on the priority tree by selecting the child node whose trajectories have a lower makespan in the next iteration, thus exploring the space of all priority orderings.
PBS is incomplete in general; however, with a complete low-level planner, it is complete for well-formed instances of various MAPF problems [60]. Specifically, PBS with a complete low-level planner is complete for well-formed Deconflicting MCPP instances, defined as follows.
Definition 2 (Well-Formedness).
A Deconflicted MCPP instance with a given input solution $\Pi$ of MCPP instance $(G,I,R)$ is well-formed if, for all $i\in I$, path $\pi_{i}\in\Pi$ does not traverse any root vertex $r_{j}\in R$ of robot $j\neq i$.
Since the given solution $\Pi$ might not satisfy the well-formed condition, before deconflicting the given $\Pi$, we preprocess each $\pi_{i}\in\Pi$ to remove any $v\in R\setminus\{r_{i}\}$ from $\pi_{i}$, which preserves complete coverage while ensuring the completeness of PBS when, for all $i$, $G$ remains connected with all vertices $v\in R\setminus\{r_{i}\}$ removed. A sufficient condition is that $G$ remains connected with all vertices $v\in R$ removed, which is used to generate instances for our experiments in Sec. VII.
@@
Alternatively, Conflict-based Search (CBS) [61] can serve as the high-level planner. CBS is optimal for MAPF problems when an optimal low-level planner is used. However, CBS generally incurs significantly longer runtimes than PBS due to its Best-First Search on a tree whose size grows exponentially with the number of conflicts. This runtime increase is even more pronounced in our multi-goal MAPF setting, where an optimal low-level planner is notably more computationally intensive than in standard single-goal MAPF. According to [43] which applies CBS to OMG-MAPF, a special case of our Deconflicting MCPP problem with unweighted edges, CBS only scales to sequences of up to 5 goals on relatively small graphs.
@@ VI-C The Low-level Planner
The lower-level computes a trajectory $\tau_{i}$ for robot $i$ that sequentially visits vertices along the given $\pi_{i}$. It uses a reservation table that records the reserved time intervals for each vertex, ensuring that the resulting trajectory $\tau_{i}$ avoids conflicts with higher-priority trajectories.
Reservation Table: To accommodate planning in an edge-weighted graph, we adopt the idea from Safe Interval Path Planning (SIPP) [62, 63] to reserve continuous-time intervals, instead of discrete time steps in the classic vertex-time A* [64]. The reservation table is indexed by a vertex, with the corresponding entry recording the reserved intervals during which the vertex is occupied. Specifically, following the conflict definition (Definition 1), given state $(v,t)$, the reservation table records the reserved interval $[t_{p},t_{s})$ for vertex $v$, where $(v_{p},t_{p})$ and $(v_{s},t_{s})$ are the preceding and succeeding states of state $(v,t)$, respectively. Notably, the reservation also records $[0,t_{s})$ for the first state $(r_{j},0)$ and $[t_{p},+\infty)$ for the last state in each higher-priority trajectory $\tau_{j}$. The safe intervals for each vertex $v$ are obtained as the complements of its reserved intervals with respect to interval $[0,+\infty)$.
SIPP Node Expansion: The low-level planner uses A* search to generate a search tree, where each node $n$ is uniquely identified by a pair $\langle v,[lb,ub)\rangle$, consisting of a vertex $v$ and one of its safe intervals $[lb,ub)$. The $g$-value $g(n)\in[lb,ub)$ of a node $n$ stores the earliest arrival time at which the robot can stay at vertex $v$ within the safe interval $[lb,ub)$ along any path. When expanding node $n=\langle v,[lb,ub)\rangle$, the search generates a child node $n^{\prime}=\langle u,[lb^{\prime},ub^{\prime})\rangle$ with $g(n^{\prime})$ for each vertex $u$ adjacent to $v$ and each safe interval $[lb^{\prime},ub^{\prime})$ of $u$, corresponding to the action of waiting at $v$ for the minimal duration $g(n^{\prime})-w_{(v,u)}-g(n)\geq 0$, then departing from $v$ at time $g(n^{\prime})-w_{(v,u)}$ and arriving at $u$ at the earliest available time $g(n^{\prime})$. The search discards node $n^{\prime}$ if no valid arrival time $g(n^{\prime})$ is available such that $g(n^{\prime})\geq lb^{\prime}$ with the corresponding departure time $g(n^{\prime})-w_{(v,u)}<ub$. If the search generates a duplicate copy of an already generated node, it keeps only the dominating one with the smallest $g$-value.
Conflict-Free Trajectory Reconstruction: The trajectory $\tau_{i}$ for $\pi_{i}$ is generated by recursively backtracking the search tree that starts from $\langle\pi_{i}[|\pi_{i}|],\cdot\rangle$ and ends at $\langle\pi_{i}[1],\cdot\rangle$. For each SIPP node $n=\langle v,\cdot\rangle$ along the search tree, we recover a state $\mathbf{x}=(v,g(n))$ and append it to $\tau_{i}$.
Given multiple ordered goals along a path $\pi$, we design three low-level planner approaches, chaining, multi-label, and adaptive.
Chaining Approach (Incomplete): This approach iteratively computes a trajectory by constructing (trajectory) segments between consecutive vertices in $\pi$. Starting from vertex $\pi[1]$ with time $t_{1}=0$, each $j-th$ iteration (for $j=1,\ldots,|\pi|-1$) computes a segment from vertex $\pi[j]$ with the given arrival time $t_{j}$ to vertex $\pi[j+1]$ with a computed arrival time $t_{j+1}$. Each segment computation uses the aforementioned A* search, starting from the node with vertex $\pi[j]$ and a safe interval containing $t_{j}$. The computation succeeds if the search expands a goal node $n=\langle\pi[j+1],\cdot\rangle$ with arrival time $t_{j+1}=g(n)$; otherwise, it fails. The search uses the precomputed shortest-path distance from any vertex $v$ to $\pi[j+1]$ as the heuristic value for all nodes associated with $v$. The final trajectory, if all segments are successfully computed, is constructed by concatenating the segments. However, this chaining approach is incomplete [43] because it does not allow backtracking. For example, a successful trajectory may require searching from alternate safe intervals of $\pi[j]$ rather than the one containing $t_{j}$, which this approach does not account for.
Multi-Label Approach (Complete and Optimal; Novel for Continuous Time): This approach extends Multi-Label A* [65, 66, 67], originally developed for unweighted graphs, to plan a time-minimal trajectory through ordered goals $\pi[j]$ ($j=1,\ldots,|\pi|$) in a single search. This approach adds a label dimension to the state space and thus searches with nodes uniquely identified by a tuple $\langle v,[lb,ub),l\rangle$, consisting of a vertex $v$, one of the safe intervals $[lb,ub)$ of $v$, and a label $l$ indicating the current goal $\pi[l]$—resulting in a novel path-planning algorithm named Multi-Label SIPP (ML-SIPP). The search of ML-SIPP follows the same node expansion rule as standard SIPP, letting a child node inherit its label $l$ from its parent, except when its vertex is $\pi[l]$, in which case its label changes to $l+1$. The search terminates when it expands a (final-)goal node $\langle\pi[|\pi|],\cdot,|\pi|+1\rangle$. ML-SIPP uses the precomputed concatenated shortest-path distance heuristic:
$h((v,\cdot,l))=\text{dist}(v,\pi[l])+\sum_{j=l}^{|\pi|-1}\text{dist}(\pi[j],\pi[j+1]).$
Since this heuristic is admissible—never overestimating the cost of reaching the (final-)goal node—ML-SIPP is a complete and optimal low-level planner.
Adaptive Approach (Complete; Novel): The adaptive approach, shown in Alg. 5, seeks to enhance practicability by combining the efficiency of the chaining approach and the completeness of ML-SIPP. It sequentially applies ML-SIPP to subsequences of goals in path $\pi$, each time starting from a subsequence of a single goal [Line 5] and backtracking to include the previous goal whenever ML-SIPP fails to find a trajectory for the current subsequence [Lines 5-5]. The backtracking depth is limited by a small integer $b_{\text{max}}\in\mathbb{Z}^{+}$ (we set it as $5$ in Sec. VII). Upon successfully planning for a single goal, the approach resets current backtracking depth $b$ [Lines 5-5]. Our adaptive approach is complete since it calls ML-SIPP for the entire sequence of goals in $\pi$ (i.e., the multi-label approach) in the worst case when ML-SIPP fails with the window size $b_{\text{max}}$ [Lines 5-5]. In practice, we postpone the ML-SIPP call on the entire goal sequence and mark the PBS node $N$ for the multi-label approach. If the high-level PBS tree reports failure once its search stack is empty, we pick one postponed node, unmark it, and call the multi-label approach for the node with children nodes branched. This strategy greatly improves the overall efficiency of the deconflicting procedure, as each call of the multi-label approach is highly time-consuming, especially when dealing with large-scale instances.
Input: path $\pi=(v_{1},\ldots,v_{|\pi|})$ for robot $i$,  reservation table $rt$ w.r.t. $\boldsymbol{\prec}_{N}$ of PBS node $N$